RC meeting and Run Plan, Friday December 9th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Current RC: Eric King [ericking (@jlab.org or @temple.edu)]
Time: Daily at 3:30PM
Location: Counting House Conference Room

Remote Connection Details

Shift Schedule

This RC Week

  • Next empty shift is a Target Operator DAY shift on Sunday 12/11.

Upcoming RC Week

  • There are two (2) other remaining shifts to be filled before winter shutdown; I'm sure Andrew would like them filled asap so he doesn't have to chase down any shift workers.

Accelerator Status

  • Everything appears to be running smoothly at the moment.
  • Hall C is down for about a week due to a vacuum break.

Current Status

We took data on 3-He during SWING with no huge disruptions.

Møller Measurement

Møller measurement was scheduled during Friday DAY. Target holding field magnet was warm and we backed out of the Moller configuration to return to SBS 3-He production.

  • Escorted access made at 12:55 by Ellen to get an initial diagnosis.

Target Polarization

Target polarization has been good. We look to be at the ~41% value suggested by Gordon


I would estimate that we lost 2.5-3.0 hours of data taking to DAQ issues. Shift crew should make an effort to document these issues clearly.

  • (6:00PM) I received a phone call from Anu Thursday afternoon that they were having difficulties with the DAQ. This was ~45 minutes worth of time.
    • I instructed Anu to contact Zeke who was also unable to diagnose and remediate the problem.
    • I contacted Alexander who made his way in to the Lab; before he arrived at the Lab the problem resolved itself and the diagnosis remains unclear. [4098814]
  • (11:45PM) I talked to Nilanga when I came in for the crew switch around 11:45pm and he said that he had a minor problem with the DAQ around 11:15pm. This may be a more routine issue. He noted it in the shift log that Anu began [4098734]
  • (2:00AM) I received a call from the shift leader that they were having difficulty with the DAQ. There were warnings and errors about ROC1. I restarted the DAQ with no issues. [4098940]
  • (5:30AM) DAQ Problems re-erupted. [4098987] [4098994] [4099003]
    • Alexander was notified and got the DAQ working again with his HALOG entry at 6:20AM. [4099013] [4099023]

Issues Resolved

Issues remaining

  • BB GEM APV unreachable and cannot be repaired at this time.

Short Term Run Plan

  • Continue production run on He3 target, 36 uA
  • Use GEnII-noSBSGEMs DAQ config
  • Use ps3=0, ps5=6, others =-1
  • BBCal threshold at -366 mV, HCal threshold at -19 mV
  • NMR measurement every 3 - 4 hours, close to 4 hours is better

Planned Accesses to the Hall

  • I presume that repairs for the BB GEM APV will wait until shutdown.

Special Requests

Please email me ericking (@jlab.org or @temple.edu) if there's something to be added.

  • Gordon said that he would like to discuss with the other spokespersons an increase in the current being placed on the target, I informed him to start that discussion and keep me in the loop.

Shutdown Planning

  • Arun has agreed that he will begin an email chain with RadCon regarding what materials are safe/unsafe in the hall.

Notices in Effect for Shift Crew

  • Power down SBS, BB, and Detector HV before making an access to the Hall

Data Collection Status

  • 75 out of 141 days (~53.2%)
  • (NEEDS TO BE UPDATED) Total uncorrected accumulated charge on He3: 47.020 C out of 232C (20.26%)
    • 2-Pass: 13.467 C
    • (NEEDS TO BE UPDATED) 3-Pass: 33.553 C (at end of OWL)